# The Intermediate Value Theorem for Linear Transformations

### v1

@@ 3. Linear Transformations: Answers
Assume first that $m\geqslant n$. Since $A$ is of full rank, we have that the rank of $A$ is $n$. This means that $A$ is injective. Let $\mathbf{e}_{k}\in\mathbb{R}^{n}$ be the $k$-th canonical vector $\mathbf{e}_{k}=(0,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$, with the number $1$ in the $k$-th coordinate. We claim that, for each $k=1,2,\dots n$, the equation $A^{\mathsf{T}}\mathbf{x}=\mathbf{e}_{k}$ has a solution $\mathbf{x}^{k}\geqslant\mathbf{0}$, $\mathbf{x}^{k}\in\mathbb{R}^{m}$; for it not, Farkas’ Lemma would imply that there exist $\mathbf{y}\in\mathbb{R}^{n}$ such that $\mathbf{y}^{\mathsf{T}}A^{\mathsf{T}}\geqslant 0$ and $\mathbf{y}^{\mathsf{T}}\mathbf{e}_{k}<0$. But this is equivalent to saying that $A\mathbf{y}\geqslant 0$ and the $k$-th entry of $\mathbf{y}$ is negative. But since $A$ is injective, if we set $\mathbf{b}:=A\mathbf{y}$, the only solution to the equation $A\mathbf{x}=\mathbf{b}$ is $\mathbf{x}=\mathbf{y}$. Since the $k$-th entry of $\mathbf{y}$ is negative, this contradicts the hypothesis. Therefore, for each $k$ we have that there exists $\mathbf{x}^{k}\geqslant\mathbf{0}$ such that
@@
There is a case in which the above two propositions can be combined into an easy necessary and sufficient condition as the next corollary states.
@@
Observe that the corollary above implies that the answer to Question 4 is affirmative, for $m=1$, if and only if $A$ has a positive entry. This has a nice geometric interpretation: if we think of the matrix $A$ as a vector in $\mathbb{R}^{n}$, in order to have an $n\times 1$ matrix $B$ with $AB=1$ and with $B\geqslant 0$, it is necessary and sufficient that the dot product of the “vector” $A$ with a vector in the first orthant [^5] The first orthant of $\mathbb{R}^{n}$ is the set of vectors $\mathbf{x}\in\mathbb{R}^{n}$ with $\mathbf{x}\geqslant\mathbf{0}$ of $\mathbb{R}^{n}$ be positive.; i.e., its angle with a vector in the first orthant of $\mathbb{R}^{n}$ is less than $90$ degrees: this is possible as long as the “vector” $A$ is not in the opposite orthant of the first orthant; i.e., at least one entry of $A$ is positive (compare with [6, Theorem 1]).
@@
Now, to finish this section, let us briefly explore what can we do to answer Question 4 if the matrix $A$ is not of full rank. So assume that we are given an $m\times n$ matrix $A$ and a vector $\mathbf{b}\geqslant\mathbf{0}$ such that the equation $A\mathbf{x}=\mathbf{b}$ has a solution and we want to find $\mathbf{x}^{\prime}\geqslant 0$ such that $A\mathbf{x}^{\prime}=\mathbf{b}$. Of course, the natural course of action is to find all solutions to $A\mathbf{x}=\mathbf{b}$ and check if we can find a nonnegative one. The usual procedure is to find an invertible $m\times m$ matrix $Q$ such that
@@
where $P$ is an $n\times n$ permutation matrix, $I_{k}$ is the $k\times k$ identity matrix and $S$ is a $k\times(n-k)$ matrix, and the zero matrices, denoted by 00, are of apropriate sizes (here $k$ is the rank of $A$, of course). We should point out that in some cases, there may be several choices of matrices $Q$ and $P$ (and $S$).
@@
and observe that $A(-3,0,1)^{\mathsf{T}}=(1,0,0)^{\mathsf{T}}=:\mathbf{b}$. Then, we can find, by elementary row operations, the matrix
@@
4. Final remarks
To summarize what we have done in this article, if $A$ is a weakly monotone matrix, the answer to question 4 is affirmative. Of course, as we mentioned before, characterizing which matrices are weakly monotone is not a simple problem (again, see [8]). Nevertheless, checking if a single matrix $A$ is weakly monotone or not is a straightforward (but perhaps not computationally easy) problem, as we mentioned at the end of the previous section.
Surprinsingly perhaps, the answer to Question 3 is easier. The $m\times n$ matrix $A$ is monotone if and only if it has a right inverse (and hence $m\geqslant n$), and if this occurs then the answer to Question 3 is affirmative. Of course, as we mentioned before, fewer matrices will have this stronger property.
The attentive reader will have noticed that Questions 3 and 4 could be rephrased by changing the direction of the inequalities in the domain (or in the range). But clearly this problem can be solved by considering the matrix $-A$ instead of the matrix $A$.
We should point out that there exists other generalizations of the Intermediate Value Theorem in several dimensions; e.g. the Poincaré-Miranda theorem (see, for example, [4]).
Lastly, we remark that there are several applications of the study of monotone matrices, weakly monotone matrices (and other types of monotonicities). Indeed, the concept of monotonicity comes from the study of error analysis and the convergence of iterative methods in numerical analysis, and has several applications in linear programming. See [2, 1, 7] and the references therein for more information.

### v2

@@ Abstract.
Assume first that $m\geqslant n$. Since $A$ is of full rank, we have that the rank of $A$ is $n$. This means that $A$ is injective. Let $\mathbf{e}_{k}\in\mathbb{R}^{n}$ be the $k$-th canonical vector $\mathbf{e}_{k}=(0,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$, with the number $1$ in the $k$-th coordinate. We claim that, for each $k=1,2,\dots n$, the equation $A^{\mathsf{T}}\mathbf{x}=\mathbf{e}_{k}$ has a solution $\mathbf{x}^{k}\geqslant\mathbf{0}$, $\mathbf{x}^{k}\in\mathbb{R}^{m}$; for if not, Farkas’ Lemma would imply that there exists $\mathbf{y}\in\mathbb{R}^{n}$ such that $\mathbf{y}^{\mathsf{T}}A^{\mathsf{T}}\geqslant 0$ and $\mathbf{y}^{\mathsf{T}}\mathbf{e}_{k}<0$. But this is equivalent to saying that $A\mathbf{y}\geqslant 0$ and the $k$-th entry of $\mathbf{y}$ is negative. But since $A$ is injective, if we set $\mathbf{b}:=A\mathbf{y}$, the only solution to the equation $A\mathbf{x}=\mathbf{b}$ is $\mathbf{x}=\mathbf{y}$. Since the $k$-th entry of $\mathbf{y}$ is negative, this contradicts the hypothesis. Therefore, for each $k$ we have that there exists $\mathbf{x}^{k}\geqslant\mathbf{0}$ such that
@@
There is a case in which Propositions 5 and 6 can be combined into an easy necessary and sufficient condition as the next corollary states.
@@
Observe that the corollary above implies that the answer to Question 4 is affirmative, for $m=1$, if and only if $A$ has a positive entry. This has a nice geometric interpretation: if we think of the matrix $A$ as a vector in $\mathbb{R}^{n}$, in order to have an $n\times 1$ matrix $B$ with $AB=1$ and with $B\geqslant 0$, it is necessary and sufficient that the dot product of the “vector” $A$ with a vector in the first orthant [^5] The first orthant of $\mathbb{R}^{n}$ is the set of vectors $\mathbf{x}\in\mathbb{R}^{n}$ with $\mathbf{x}\geqslant\mathbf{0}$. of $\mathbb{R}^{n}$ be positive; i.e., its angle with a vector in the first orthant of $\mathbb{R}^{n}$ is less than $90$ degrees. This is possible as long as the “vector” $A$ is not in the opposite orthant of the first orthant; i.e., at least one entry of $A$ is positive (compare with [8, Theorem 1]).
@@
Now, to finish this section, let us briefly explore what can we do to answer Question 4 if the matrix $A$ is not of full rank. So assume that we are given an $m\times n$ matrix $A$ and a vector $\mathbf{b}\geqslant\mathbf{0}$ such that the equation $A\mathbf{x}=\mathbf{b}$ has a solution and we want to find $\mathbf{x}^{\prime}\geqslant 0$ such that $A\mathbf{x}^{\prime}=\mathbf{b}$. The natural course of action is to find all solutions to $A\mathbf{x}=\mathbf{b}$ and check if we can find a nonnegative one. The usual procedure is to find an invertible $m\times m$ matrix $Q$ such that
@@
where $P$ is an $n\times n$ permutation matrix, $I_{k}$ is the $k\times k$ identity matrix and $S$ is a $k\times(n-k)$ matrix, and the zero matrices, denoted by 00, are of appropriate sizes (here $k$ is the rank of $A$, of course). We should point out that in some cases, there may be several choices of matrices $Q$ and $P$ (and $S$).
@@
this leads to $A(-3,0,1)^{\mathsf{T}}=(1,0,0)^{\mathsf{T}}=:\mathbf{b}$. Then, we can find, by elementary row operations, the matrix
@@
4. Final remarks.
To summarize what we have done in this article, if $A$ is a weakly monotone matrix, the answer to Question 4 is affirmative. As we mentioned before, characterizing which matrices are weakly monotone is not a simple problem (again, see [10]). Nevertheless, checking if a single matrix $A$ is weakly monotone or not is a straightforward (but perhaps not computationally easy) problem, as we mentioned in the previous section.
Surprinsingly perhaps, the answer to Question 3 is easier. The $m\times n$ matrix $A$ is monotone if and only if it has a left inverse (and hence $m\geqslant n$), and if this occurs then the answer to Question 3 is affirmative. Fewer matrices will have this stronger property.
We should mention that the answers to Questions 3 and 4 above have important applications. Collatz [3] introduced the concept of mononote matrices in his study of the approximation of solutions of systems of linear equations. Indeed, for a nonsingular square matrix $A$, if we need to solve the equation $A\mathbf{x}=\mathbf{b}$, and we have found vectors $\mathbf{v}_{1}\leqslant\mathbf{v}_{2}$ such that $A\mathbf{v}_{1}\leqslant\mathbf{b}\leqslant A\mathbf{v}_{2}$, it seems natural to expect that the solution $\mathbf{x}$ might be found “between” $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$; i.e., $\mathbf{v}_{1}\leqslant\mathbf{x}\leqslant\mathbf{v}_{2}$. Collatz characterized those matrices for which this occurs: monotone matrices.
It is not surprising then than the concept of monotonicity is crucial in the study of iterative methods for solving systems of linear equations. Richard S. Varga [11] introduced the concept of regular splitting of a matrix $A$, which includes the monotonicity of the matrix $A$, to find necessary conditions for the convergence of certain iterative methods for approximating the solutions to $A\mathbf{x}=\mathbf{b}$. See [2, 1, 9] and the references therein for more information on these topics, including some applications to linear programming.
What is it about linear maps that allowed us to partially solve Question 4? Observe that in Proposition 5, in the case $m\geqslant n$ we used the existence of a left inverse that sends nonnegative vectors to nonnegative vectors. In general, for a function $f:\mathbb{R}^{m}\to\mathbb{R}^{n}$, if there exists a function $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ which sends nonnegative vectors to nonnegative vectors, and such that $g\circ f$ is the identity on $\mathbb{R}^{m}$, then we would have that if $\mathbf{0}\leqslant f(\mathbf{x})$, then $\mathbf{0}\leqslant g(f(\mathbf{x}))=\mathbf{x}$, so condition (2) of Proposition 1 is satisfied (of course, since $f$ is not necessarily linear, we do not have the equivalence in Proposition 1). When is there a left inverse $g$ to a function $f$, such that $g$ maps the first orthant into itself? Perhaps nonlinear versions of Farkas’ Lemma (e.g, [4, 12, 5]) might give a partial answer. We leave this question open for the interested reader.
Lastly, we should point out that there exist other generalizations of the Intermediate Value Theorem in several dimensions; e.g., the Poincaré-Miranda theorem (see, for example, [6]).
Acknowledgments. The author wishes to thank Levent Ulku, who approached the author with a question that led to the topic of this paper. Also, the author thanks Blair Madore, Edgar Possani and the anonymous reviewers for their comments on the first version of this manuscript, which significantly improved the exposition.
This work is partially supported by the Asociación Mexicana de Cultura A.C.
2020 Mathematics Subject Classification: Primary 15A24, Secondary 15A39, 15B48.